# Stealing Trust: Unraveling Blind Message Attacks in Web3 Authentication

### v1

@@ 4. Blind Message Attack
This section presents Blind Message Attacks, which exploit vulnerabilities in the Web3 authentication implementation as described in Section 3 to gain unauthorized access. Specifically, when a user logs into a malicious website using Web3 authentication, it tricks the user into signing a message from another website, thereby gaining access to the user’s account on that other website. This section first defines the threat model, followed by a motivating example. Then, we analyzed the potential vulnerabilities in the Web3 authentication process. Finally, according to the impact of these vulnerabilities, we classify the risk levels of Blind Message Attacks and discuss advanced attacks. Attack demos are available at https://sites.google.com/view/web3auth.
@@
In our threat model, the attacker’s objective is to steal the identities of users on target websites, thereby gaining unauthorized access. This occurs when a user logs into the attacker’s (malicious) website using Web3 authentication. The attacker exploits this opportunity to steal the user’s identity (signature) of the target website. In our model, the attacker is not capable of exploiting browser vulnerabilities or cracking users’ crypto wallets. Also, attackers cannot intercept and decrypt network packets, such as man-in-the-middle (MitM) attacks. The attacker’s capabilities are limited to exploiting vulnerabilities in the Web3 authentication process of the target website.
@@ 4.1. Motivating Example
In Steps 1-2, a user visits a malicious website and connects to the wallet. The malicious website queries the transaction record of the address (public key) in the blockchain, identifies the websites that the user has logged into before by contract addresses, and then selects the most valuable target website based on information such as balance, trading volume.
@@
In Steps 11-14, the malicious website uses the signature, message, and address previously obtained to request a token from the target website, and then holds the token to access the user’s account.
Figure 2. Blind Message Attack.
@@ 4.3. Security Risks
This section discusses the impact of vulnerabilities on attacks and discusses advanced attacks, namely Replay Attacks and Blind Multi-Message Attacks.
@@ 4.3.1. Risk Levels
Based on the vulnerability’s applicability to Blind Message Attacks in Section 4.2, we categorize the attacks into four levels. In critical to medium-risk scenarios, users are unable or difficult to detect the attack. Therefore, we consider there to be significant security issues in the implementation of Web3 authentication on these websites. In low-risk scenarios, attentive users might identify the message as originating from another website, thereby preventing the attack.
@@ 4.3.2. Advanced Attacks
Generally, an authentication token has a specified expiration time, and once expired, the attacker cannot access the user’s account without securing a new signature. However, if the message lacks nonce (V1) or the verification process is flawed (V2, V3), an attacker can replay the signature repeatedly to refresh the token, thus maintaining unauthorized access. It is important to note that while broader replay attacks could be carried out by intercepting signatures, such as man-in-the-middle attacks, such cases fall outside the threat model this paper considers. Our focus remains on assessing the impact of replay attacks within the framework of blind message attacks.
@@
Attackers can exploit vulnerabilities across multiple websites to create a crafted message that simultaneously passes Web3 authentication on all these websites. As a result, with a single signature from the user, the attacker can access the user’s accounts on several websites. The crafted message is equivalent to multiple messages on websites, so we call the attack Blind Multi-Message Attack. Specifically, for websites that do not verify the message body (V2) or only verify its presence rather than strict equality (V3), an attacker can bypass the server verification of these websites by constructing a message that contains the required fields. We will describe this attack in detail with a specific case in Section 6.2.
@@ 5.1. Overview
Our tool, Web3AuthChecker, interfaces with these requests to identify vulnerabilities. In simple terms, it fills different attack payloads into HTTP requests, then analyzes whether responses are as expected. The architecture, depicted in Figure 3, is composed of Checker and FlexRequest. There are some predefined parsers in FlexRequest to support converting different types of API and configuration files into a unified Request items object. The vulnerability detection process is as follows:
@@
    FlexRequest parses the website’s response according to the pre-defined configuration and returns the parsed results back to Checker.
@@ 5.2. Checker Implementation
Message Checker begins by sending a QUERY request to retrieve a message from the website. Then, it matches the presence of domain names (domain) and website names (name). To detect vulnerabilities in verification, the checker modifies the message to a random string as the attack payload and signs it. If a subsequent AUTH request returns a token, it indicates that the server does not adequately check the message. To further verify, the checker 1) empties the message body and 2) appends random characters at the beginning of the message, to assess any vulnerabilities in the server’s message body verification. If the message fails any of the above checks, it is considered vulnerable to Blind Message Attacks, and the checker returns specific security risk levels based on the type of vulnerability.
@@
    It generates a value similar to the nonce and replaces it, before sending the fourth AUTH.
@@ 5.3. FlexRequest: A Python HTTP Library
FlexRequest supports using keys to replace dynamic parameters in the API. Before executing a request, it automatically substitutes these keys with the corresponding values. After receiving the response, FlexRequest retrieves the values at the specified positions according to a predefined configuration, binding them to the corresponding keys. By managing the values of these keys, developers can perform unified testing across multiple APIs, regardless of their differences.
Furthermore, FlexRequest maintains a session context throughout a session to pass the previous response values between a series of requests. For example, the message obtained from the Query response will be stored in the session context and bound to the key msg. When conducting an Auth request, FlexRequest automatically populates the request with the value of msg from the session context. For detailed insights into API testing challenges and FlexRequest’s unique solutions, see Appendix B.
@@
colsep=2pt,rowsep=0pt,stretch = 1 {tblr} width = 1colspec =X[0.5,c]X[2.2,c]X[1.5,c]X[0.15,c] X[0.9,c]X[0.7,c]X[0.7,c]X[0.22,c] X[1,c]X[0.7,c]X[0.8,c]X[1.1,c]X[1,c]X[0.22,c] X[0.7,c]X[1,c]X[0.8,c], rowodd = white, roweven = gray!25, row1 = white, font=, row2 = white, font=, column1 = colsep=4pt, font=, cell11 = r=2c, cell12 = r=2c, cell13 = r=2c, cell15 = c=3c, cell19 = c=5c, cell115 = c=3c, # Website Category Message Design (V1) Server Verification Security Risk
@@
In the Web3 authentication process, cases 1-25 are for user login, while cases 26-29 are for profile update.
@@ 6. Findings
Given the limited research on Web3 authentication, there are no existing relevant datasets. DappRadar, a well-known Web3 dapp distribution platform, provided a basis for our collection: we selected 18 marketplaces that support Web3 authentication from the top 50 in DappRadar’s Top Decentralized Marketplaces list (dap, 2024). Additionally, we identified nine websites that support Web3 authentication through Google searches. Of these 27 websites, two required Web3 authentication for both user login and profile updates (4 test cases). Among the remaining 25 websites, 23 utilized Web3 authentication solely for user login (23 test cases), and two employed it specifically for profile updates (2 test cases). Therefore, we have a total of 29 test cases. According to DappRadar’s statistics, in January 2024 alone, the total transaction volume of 16 of these websites exceeded 592 million US dollars, and the number of unique active wallets (UAW) surpassed 1.29 million.
@@
We thoroughly examined each website and collected the relevant APIs, exporting them as JSON files in Postman format. We also prepared corresponding configuration files for Web3AuthChecker. We deployed Web3AuthChecker on Github Codespaces, configured with 4 cores and 8GB RAM. The curl-impersonate browser was set to Chrome (v110), and the API response timeout was limited to 10 seconds. To avoid being blocked due to frequent requests, we instituted a one-minute interval between each request.
@@ 6.1. Analysis of Results
Table 5.3 provides detailed information and detection results for all 29 test cases. Web3AuthChecker examines whether messages include domain, name, and nonce, and it verifies if the server checks the message, message body (body), nonce, signature, and address to identify the risks defined in Section 4. The results are alarming, with 22 out of the 29 test cases found to be subject to medium to critical risks making them vulnerable to Blind Message Attacks. At the same time, we identified critical risks on two cases, and we have informed the respective vendors. One of them has already acknowledged the risk and fixed it. Also, 11 cases are at the risk of Replay Attacks, and seven have Blind Multi-Message Attacks. From Table 5.3, it can be observed that each website has its own distinct vulnerabilities, indicating that their Web3 authentication implementations are independent. Therefore, we conclude that there are no common server-side SDKs across these websites.
@@
In nine high-risk cases, either the domain and name fields do not exist in the message (V1), or the server does not verify message body, so the attacker can tamper with them (V2). The nine high-risk websites provide attractive targets for attackers. By exploiting the users’ inability to identify the source of messages, malicious websites can effortlessly trick users into signing messages from these websites. Notably, in January 2024, Apeiron, QuestN, and MOBOX, all identified as high-risk websites, reported 89.21k, 54.94k, and 27.26k Unique Active Wallets (UAW) respectively. These UAWs underscore the substantial user engagement and the potential impact of such vulnerabilities.
In the seven medium-risk cases, the attacker may need more sophisticated techniques to implement the attack, such as using a domain name identical to the target website (V1) or hiding the target website’s message body within the malicious website’s message (V3). As discussed in the case study section (Section 6.2), these hurdles may not be too challenging for experienced attackers. Importantly, even cases categorized as low risk are not necessarily exempt from blind message attacks, as attackers still have opportunities to deceive unwary users.
Vulnerabilities in verification can weaken a well-structured design. For instance, Element (#5), Galler (#12), and Planetix (#21) all have appropriately structured messages containing the essential fields. However, Element does not improperly verify the message body (V3), thus placing it at medium risk, while the other two do not verify the message body (V2), which puts them at high risk. Notably, as of January 2024, Element reported 158.63k UAWs and a transaction volume of \$18.91M. These security risks emphasize the importance of server verification.
@@ 6.2. Case Study
LearnBlockchain is a well-known blockchain community with more than 2 million users. It supports both traditional password-based authentication and Web3 authentication. In the Web3 authentication, LearnBlockchain’s message is simply learnblockchain. Moreover, the server only verifies the signature and does not perform any checks on the message, allowing any signature of the user to pass the Web3 authentication. This vulnerability provides an easy way for an attacker to gain unauthorized access to a user account, which can lead to the exposure of important personal information and potential asset losses. Specifically, LearnBlockchain’s points system supports points exchange for currency, so attackers can transfer points from the victim’s account to their own account and then withdraw them.
We promptly reported blind message attack and replay attack to LearnBlockchain, who fixed their Web3 authentication based on our suggestions and awarded us 2500 points for our responsible disclosure. Another website, Galler, facing similar security issues, was also notified about the vulnerabilities. However, at the time of writing, we have not received any response from Galler.
@@ 7. Mitigation: Web3AuthGuard
Given the vulnerabilities identified by Web3AuthChecker, extensive server-side updates are necessary but challenging to implement quickly. This section introduces Web3AuthGuard, a solution designed to detect potential attacks on the user (wallet) side. Compared to Web3AuthChecker, Web3AuthGuard automatically detects potential attacks during wallet operation (signature request), without the need for pre-configuration.
@@ 7.1. Design of Web3AuthGuard
Web3AuthGuard is designed to detect suspicious messages in crypto wallets, aiming to prevent blind message attacks during the Web3 authentication process. The workflow of Web3AuthGuard is depicted in Figure 4. During the Web3 authentication process, a website triggers the wallet’s signature API (Personal Sign). Consequently, the wallet prompts a Signature Request page for the user to sign the message. For wallets featuring Web3AuthGuard, the first step involves comparing the new message with previously signed messages (message templates). If the new message significantly resembles a previously signed message from a different website, Web3AuthGuard promptly raises an alert on the top of the signature request page. This alert notifies the user of a potential Blind Message Attack, enabling them to make an informed decision about whether to proceed with the authentication process. The above process mainly relies on three components of Web3AuthGuard: template extraction, fuzzy matching and wallet alerts. Below we introduce them in detail.
@@
The template extraction process is as follows: when the user logs into the website for the first time, the wallet will record the message and mark it with the domain name. When the user logs in to the website again, Web3AuthGuard will compare the new message with the previously stored message to extract the message template. Specifically, the process involves a word-by-word comparison of the two messages. Words that match are kept and incorporated into the template. In contrast, non-matching words, pinpointed by predefined variable fields, are substituted with corresponding wildcards in the template. This updated template then supersedes the previously saved message and becomes the reference for detecting blind message attacks.
@@
Fuzzy Matching
Web3AuthGuard employs regular expression matching to compare the new message with the stored message template during the Web authentication process. If a match is found, indicating similarity between the new message and the template, Web3AuthGuard further checks if they originate from different websites. If they do, Web3AuthGuard triggers an alert on the signature request page of the wallet. This alert warns users of a potential blind message attack, indicating that the website may be malicious.
@@
Web3AuthGuard will add alerts to the signature request page to inform users of potential security risks. There are two types of alerts, blind message attack and lack of domain name, which are independent and may occur at the same time. If fuzzy matching detects a blind message attack, Web3AuthGuard will add a red alert, as shown in Figure 5(a). The alert will prompt for the domain names of the current website and victim website. The user should check whether the domain name (domain) or the similar website name (name) of the victim website exists in the message. If it exists, it means that it is a blind message attack, and the user should click the Reject button on the signature page to refuse the signature. If the domain and name do not exist in the message, this may be a potential attack. Because once the user signs, the website will intentionally or unintentionally gain the access of the user’s account on the victim website. Therefore, we still recommend that users do not sign the message.
In addition, Web3AuthGuard will match whether the domain name (domain) of the visited website appears in the message. If it does not exist, Web3AuthGuard will raise a yellow alert (lack of domain name), as shown in the figure  5(b). We recommend that users do not sign the message as it may originate from other websites or be exploited by other malicious websites. Finally, users should report any malicious activity on the website to the wallet, and most wallets will then warn other users about these dangerous websites.

### v4

@@ 4. Blind Message Attack
This section presents Blind Message Attacks, which exploit vulnerabilities in the Web3 authentication implementation described in Section 3 to gain unauthorized access. Specifically, when a user logs into a malicious website using Web3 authentication, it tricks the user into signing a message from another website, thereby gaining access to the user’s account on that other website. This section first defines the threat model, followed by a motivating example. Then, we analyzed the potential vulnerabilities in the Web3 authentication process. Finally, according to the impact of these vulnerabilities, we classify the risk levels of Blind Message Attacks and discuss advanced attacks. Attack demos are available at https://sites.google.com/view/web3auth.
@@
In our threat model, the attacker aims to steal users’ identities on target websites, thereby gaining unauthorized access. This occurs when a user logs into the attacker’s (malicious) website using Web3 authentication. The attacker exploits this opportunity to steal the user’s identity (signature) of the target website. In our model, the attacker cannot exploit browser vulnerabilities or crack users’ crypto wallets. Also, attackers cannot intercept and decrypt network packets, such as man-in-the-middle (MitM) attacks. The attacker’s capabilities are limited to exploiting vulnerabilities in the Web3 authentication process of the target website.
Figure 2. Blind Message Attack.
@@ 4.1. Motivating Example
In Steps 1-2, a user visits a malicious website and connects to the wallet. The malicious website queries the transaction record of the address (public key) in the blockchain, identifies the websites that the user has logged into before by contract addresses, and then selects the most valuable target website based on information such as balance and trading volume.
@@
In Steps 11-14, the malicious website uses the signature, message, and address previously obtained to request a token from the target website and then holds the token to access the user’s account.
@@ 4.3. Security Risks
This section discusses the impact of vulnerabilities on Blind Message Attacks and discusses two advanced attacks, namely Replay Attacks and Blind Multi-Message Attacks.
@@ 4.3.1. Risk Levels
Based on the vulnerability analysis in Section 4.2, we categorize Blind Message Attacks into four levels. In critical to medium-risk scenarios, users are often unable to detect the attack or find it challenging to do so. In low-risk scenarios, attentive users might identify the message as originating from another website, thereby preventing the attack.
@@ 4.3.2. Advanced Attacks
Generally, an authentication token has a specified expiration time, and once it expires, the attacker cannot access the user’s account without securing a new signature. However, if the message lacks nonce (V1) or the verification process is flawed (V2, V3), an attacker can replay the signature repeatedly to refresh the token, thus maintaining unauthorized access. It is important to note that while broader replay attacks could be carried out by intercepting signatures, such as man-in-the-middle attacks, such cases fall outside the threat model this paper considers. Our focus remains on assessing the impact of replay attacks within the framework of Blind Message Attacks.
@@
Attackers can exploit vulnerabilities across multiple websites to create a crafted message that simultaneously passes Web3 authentication on all these websites. As a result, with a single signature from the user, the attacker can access the user’s accounts on several websites. The crafted message is equivalent to multiple messages on websites, so we call the attack Blind Multi-Message Attack. Specifically, for websites that do not verify the message body (V2) or only verify its presence rather than strict equality (V3), an attacker can bypass the server verification of these websites by constructing a message that contains the required fields. We will describe this attack in detail in Section 6.2.
@@ 5.1. Overview
Our tool, Web3AuthChecker, interfaces with these requests to identify vulnerabilities. In simple terms, it injects different attack payloads into HTTP requests, then analyzes whether responses are as expected. The architecture, depicted in Figure 3, is composed of Checker and FlexRequest. There are some predefined parsers in FlexRequest to support converting different types of API and configuration files into a unified Request items object. The vulnerability detection process is as follows:
@@
    FlexRequest parses the website’s response according to the predefined configuration and returns the parsed results back to Checker.
@@ 5.2. Checker Implementation
Message Checker begins by sending a QUERY request to retrieve a message from the website. Then, it matches the presence of domain names (domain) and website names (name). To detect vulnerabilities in verification, the checker modifies the message to a random string as the attack payload and signs it. If a subsequent AUTH request returns a token, it indicates that the server does not adequately check the message. To further verify, the checker 1) empties the message body and 2) appends random characters at the beginning of the message to assess any vulnerabilities in the server’s message body verification. If the message fails any of the above checks, it is considered vulnerable to Blind Message Attacks, and the checker returns specific security risk levels based on the type of vulnerability.
@@
    It generates a value similar to the nonce and replaces it before sending the fourth AUTH.
@@ 5.3. FlexRequest: A Python HTTP Library
FlexRequest supports using keys to replace dynamic parameters in the API. Before executing a request, it automatically substitutes these keys with the corresponding values. After receiving the response, FlexRequest retrieves the values at the specified positions according to a predefined configuration, binding them to the corresponding keys. By managing the values of these keys, developers can perform unified testing across various APIs.
Furthermore, FlexRequest maintains a session context throughout a session to pass the previous response values between a series of requests. For example, the message obtained from the Query response will be stored in the session context and bound to the key msg. When conducting an Auth request, FlexRequest automatically populates the request with the value of msg from the session context. For detailed insights into API testing challenges and FlexRequest’s unique solutions, please refer to our open-source tool, Web3AuthChecker  [^3] https://github.com/d0scoo1/Web3AuthChecker.
@@
colsep=2pt,rowsep=0pt,stretch = 1 {tblr} width = 1colspec =X[0.5,c]X[2,c]X[1.5,c]X[0.15,c] X[0.9,c]X[0.7,c]X[0.7,c]X[0.2,c] X[1,c]X[0.7,c]X[0.8,c]X[1.1,c]X[1,c]X[0.2,c] X[0.8,c]X[0.8,c]X[1,c], rowodd = white, roweven = gray!25, row1 = white, font=, row2 = white, font=, column1 = colsep=4pt, font=, cell11 = r=2c, cell12 = r=2c, cell13 = r=2c, cell15 = c=3c, cell19 = c=5c, cell115 = c=3c, # Website Category Message Design (V1) Server Verification Security Risk
@@
[…]
@@ 6. Findings
Given the limited research on Web3 authentication, no relevant datasets exist. DappRadar, a well-known Web3 dapp distribution platform, provided a basis for our collection: we selected 18 marketplaces that support Web3 authentication from the top 50 in DappRadar’s Top Decentralized Marketplaces list (dap, 2024). Additionally, we identified nine websites that support Web3 authentication through Google searches. Two of these 27 websites required Web3 authentication for user login and profile updates (4 test cases). Among the remaining 25 websites, 23 utilized Web3 authentication solely for user login (23 test cases), and two employed it specifically for profile updates (2 test cases). Therefore, we have a total of 29 test cases. According to DappRadar’s statistics, in January 2024 alone, the total transaction volume of 16 websites exceeded 592 million US dollars, and the number of unique active wallets (UAW) surpassed 1.29 million.
@@
We thoroughly examined each website and collected the relevant APIs, exporting them as JSON files in Postman format. We also prepared corresponding configuration files for Web3AuthChecker. We deployed Web3AuthChecker on Github Codespaces, configured with 4 cores and 8GB RAM. We set Chrome (v110) as the curl-impersonate browser and limited the API response timeout to 10 seconds. To avoid being blocked due to frequent requests, we instituted a one-minute interval between each request.
@@ 6.1. Analysis of Results
Table 5.3 provides detailed information and detection results for all 29 test cases. Web3AuthChecker examines whether messages include domain, name, and nonce, and it verifies if the server checks the message, message body (body), nonce, signature, and address to identify the risks defined in Section 4. The results are alarming, with 22 out of the 29 test cases found to be subject to medium to critical risks, making them vulnerable to Blind Message Attacks. At the same time, we identified critical risks in two cases, and we have informed the respective vendors. One of them has already acknowledged the risk and fixed it. Also, 11 cases are at the risk of Replay Attacks, and seven have Blind Multi-Message Attacks. Table 5.3 shows that each website has its own distinct vulnerabilities, indicating that their Web3 authentication implementations are independent. Therefore, we conclude that these websites have no common server-side SDKs.
@@
In nine high-risk cases, either the domain and name fields do not exist in the message (V1), or the server does not verify the message body so the attacker can tamper with them (V2). The nine high-risk websites provide attractive targets for attackers. By exploiting the users’ inability to identify the source of messages, malicious websites can effortlessly trick users into signing messages from these websites. Notably, Apeiron (#15), QuestN (#19), and MOBOX (#22), all identified as high-risk websites, reported 89.21k, 54.94k, and 27.26k Unique Active Wallets (UAW), respectively, in January 2024. These UAWs underscore the substantial user engagement and the potential impact of such vulnerabilities.
In the seven medium-risk cases, the attacker may need more sophisticated techniques to implement attacks, such as using a domain name identical to the target website (V1) or hiding the target website’s message body within the malicious website’s message (V3). As discussed in the case study section (Section 6.2), these hurdles may be relatively easy for experienced attackers. Importantly, even cases categorized as low risk are not necessarily exempt from Blind Message Attacks, as attackers still have opportunities to deceive unwary users.
Vulnerabilities in verification can weaken a well-structured design. For instance, Element (#5), Galler (#12), and Planetix (#21) all have appropriately structured messages containing the essential fields. However, Element does not improperly verify the message body (V3), thus placing it at medium risk. In contrast, the other two do not verify the message body (V2), which puts them at high risk. As of January 2024, Element reported 158.63k UAWs and a transaction volume of \$18.91M. These security risks emphasize the importance of server verification.
@@ 6.2. Case Study
LearnBlockchain (#17) is a well-known blockchain community with more than 2 million users. It supports both traditional password-based authentication and Web3 authentication. In the Web3 authentication, LearnBlockchain’s message is simply learnblockchain. Moreover, the server only verifies the signature and does not perform any checks on the message, allowing any signature of the user to pass the Web3 authentication. This vulnerability provides an easy way for an attacker to gain unauthorized access to a user account, which can lead to the exposure of important personal information and potential asset losses. Specifically, LearnBlockchain’s points system supports points exchange for currency, so attackers can transfer points from the victim’s account to their own account and then withdraw them.
We promptly reported blind message attack and replay attack to LearnBlockchain, who fixed their Web3 authentication based on our suggestions and awarded us 2500 points for our responsible disclosure. Another website, Galler (#12), facing similar security issues, was also notified about the vulnerabilities. However, at the time of writing, we have not received any response from Galler.
@@ 7. Mitigation: Web3AuthGuard
Given the vulnerabilities identified by Web3AuthChecker, extensive server-side updates are necessary but challenging to implement quickly. This section introduces Web3AuthGuard, a solution designed to detect potential attacks on the user (wallet) side. Compared to Web3AuthChecker, Web3AuthGuard automatically detects potential attacks during wallet operation (signature request) without the need for pre-configuration.
@@ 7.1. Design of Web3AuthGuard
Web3AuthGuard is designed to detect suspicious messages in crypto wallets, to prevent Blind Message Attacks during the Web3 authentication process. The workflow of Web3AuthGuard is depicted in Figure 4. During the Web3 authentication process, a website triggers the wallet’s signature API (Personal Sign). Consequently, the wallet prompts a Signature Request page for the user to sign the message. For wallets featuring Web3AuthGuard, the first step involves comparing the new message with previously signed messages (message templates). If the new message significantly resembles a previously signed message from a different website, Web3AuthGuard promptly raises an alert on the top of the signature request page. This alert notifies the user of a potential Blind Message Attack, enabling them to decide whether to proceed with the authentication process. The above process mainly relies on three components of Web3AuthGuard: template extraction, fuzzy matching, and wallet alerts. Below, we introduce them in detail.
@@
Template extraction process is as follows: when the user logs into the website for the first time, the wallet will record the message and mark it with the domain name. When the user logs in to the website again, Web3AuthGuard will compare the new message with the previously stored message to extract the message template. Specifically, the process involves a word-by-word comparison of the two messages. Words that match are kept and incorporated into the template. In contrast, non-matching words, pinpointed by predefined variable fields, are substituted with corresponding wildcards in the template. This updated template then supersedes the previously saved message and becomes the reference for detecting blind message attacks.
@@
Fuzzy Matching
Web3AuthGuard employs regular expression matching to compare the new message with the stored message template during the authentication process. If a match is found, indicating similarity between the new message and the template, Web3AuthGuard further checks if they originate from different websites. If they do, Web3AuthGuard adds an alert on the signature request page, indicating a potential blind message attack.
@@
Web3AuthGuard will add alerts to the signature request page to inform users of potential security risks. There are two types of alerts, blind message attack and lack of domain name, which are independent and may occur at the same time. If fuzzy matching detects a blind message attack, Web3AuthGuard will add a red alert, as shown in Figure 5(a). The alert will prompt for the domain names of the current and victim websites. The user should check whether the domain name (domain) or the similar website name (name) of the victim website exists in the message. If it exists, it is a blind message attack, and the user should click the Reject button on the signature page to refuse the signature. If the domain and name do not exist in the message, this may be a potential attack. Once the user signs, the website will intentionally or unintentionally gain access to the user’s account on the victim’s website. Therefore, we still recommend that users do not sign it.
In addition, Web3AuthGuard will match whether the domain of the visited website appears in the message. If it does not exist, Web3AuthGuard will raise a yellow alert (lack of domain name), as shown in the figure  5(b). We recommend that users do not sign the message as it may originate from other websites or be exploited by other malicious websites. Finally, users should report any malicious activity on the website to their wallets, and most wallets will then warn other users about these dangerous websites.